OpenStudy (anonymous):

how do i solve d/dx 20log(4x)+1000x^2

OpenStudy (amistre64):

by deriving id assume :)

OpenStudy (amistre64):

the complicated issue is in recalling how to deal with the log(x)

OpenStudy (anonymous):

agree. is it 20/x? but what happens to 4x?

OpenStudy (amistre64):

ln(x) derives down to 1/x; so lets translate that log into lns

OpenStudy (amistre64):

log(4x) = ln(4x)/ln(10) by change of base right?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=d%2Fdx20log%284x%29%2B1000x^2 click "show steps"

OpenStudy (amistre64):

ln(10) is a constant so it can be pulled out; ln(4x) becomes 1/4x and the chain rule pops out a 4 to get: 4/4x = 1/x; bring back our straglers to get: 20/(x ln(10))

OpenStudy (anonymous):

ok. now i get where i went wrong. thx.:)

OpenStudy (amistre64):

youre welcome :)